# AI4EOSC: a Federated Cloud Platform for Artificial Intelligence in Scientific Research

### 17 Mar 2026 v2

@@ 3 The AI4EOSC platform architecture
As already introduced, the AI4EOSC Platform is designed to help scientists to develop, train and deploy AI models in a user-friendly manner. By allowing transparent access to a large set of federated computing resources, the platform ensures scalability and efficiency of scientific workflows. In the following section, we describe the architecture of AI4EOSC, using the C4 Model [13] to provide a clear and structured visualization of its components and interactions.
Our architecture follows a modular design, organized into four distinct systems: AI4EOSC Development Platform, AI4EOSC AI/ML as a Service, AI4EOSC LLM services, and AI4EOSC Platform Orchestration. These components work interconnected in order to be able to deliver an specific functionality, but being integrated with the rest of the systems. This modularity is illustrated in Figure 1. The legend for all the C4 diagrams is shown in Figure 2
Figure 1: AI4EOSC architecture system landscape view (C4 notation).
@@ 3.1 AI4EOSC Development Platform system
Figure 3: AI4EOSC Development platform architecture: container view (C4 notation).
@@ 3.1.1 Workload Management System
The Workload Management System (WMS) is a federation of service providers that is used to manage the platform resources and execute the platform jobs. Given the federated nature of the platform, we have chosen a set of components that allow us to federate geographically distributed resources easily in the form of a service mesh. Our WMS comprises the usage of Consul [43] to federate the providers (following a WAN approach) and Nomad [44] to run the workloads. The current federation includes resources from over 5 different clouds located in 5 different countries: IFCA-CSIC (Spain), IISAS (Slovakia), INCD (Portugal), Tubitak (Turkey) and PNSC (Poland).
The WMS runs both user-specific jobs, and system/platform tasks that are executed independently of the users.
User jobs include both AI modules (3.1.3) and tools (3.1.4). AI modules can run in two modes: standard and batch. In standard mode, resources are allocated to users without a predefined time limit, and users access them via an Interactive Development Environment (IDE) or an API. In batch mode, resources are allocated to users for the duration strictly needed to run the training process and then freed for other users.
In addition, we have enriched the Nomad jobs with a number of sidecar tasks to enhance the user experience transparently, like connecting the job with external storage systems, downloading and syncing datasets from external repositories at launch time or delivering user notifications.
System jobs include, apart from the sidecar tasks already mentioned, Traefik [54] reverse proxies to allow users to access their containers, the deployment of the Platform API and the Platform Dashboard (3.1.2), and other internal management and house-keeping tasks that are never exposed to the users.
@@ 3.1.2 Platform API and dashboard
Built on top of the API, the AI4EOSC Dashboard is the main user entrypoint to the AI4EOSC Platform. Both PAPI and the dashboard are built with extensibility in mind, in order to be able to easily integrate new functionality without interfering with the current one.
@@
The Dashboard (Figure 4 allows users to easily explore and view the associated metadata for the different catalogs available in the system: AI modules (3.1.3), tools (3.1.4), LLM services (3.3) and other external AI catalogs (4.2.2). It also provides the ability to launch development environments or codespaces in the platform, batch executions, easy management of existing deployments, platform-wide secret management (e.g., API keys, storage integration) and resource consumption reporting, among other user-facing functionalities.
@@ 3.1.3 AI Modules
The AI modules are the main platform asset that users interact with, as they encapsulate the AI/ML models that the platform is able to manage. They are both produced and consumed by the users, although the AI4EOSC development team also maintains reference implementations of relevant AI models. They are created and stored as git repositories (currently hosted on GitHub [^2] https://github.com/ai4os-hub, but any other software forge can be used) allowing users to easily manage and upload new updates to the modules, as well as making it possible to enforce quality assurance mechanisms and MLOps practices (3.1.5) from the platform point of view. To support the development of new AI modules that comply with the AI4EOSC platform and best practices (e.g., metadata standards (4.1)), a set of pre-defined software templates is provided to users.
@@ 3.1.4 Tools
For technical reasons, some of these tools do not make sense to be deployed per-user and therefore are instead offered as platform-wide services, like the experiment tracking service (3.1.6) or the drift monitoring (3.1.6); however the integration mechanism remains the same.
@@
The AI4EOSC Platform allows users to deploy a ready-to-use development environment, referred to as Codespaces, to create their new AI modules (3.1.3). The environment can be used out-of-the-box with multiple versions of the main Deep Learning frameworks such as TensorFlow [1], PyTorch [70].
Additional frameworks are also supported by starting from provided plain Ubuntu or NVIDIA CUDA flavours, or by installing a custom software stack.
When configuring the deployment, users can choose their preferred IDE for code development. The platform currently supports both VSCode [59] (bundled with many useful extensions, and installed using code-server [22]), and JupyterLab [50]. All IDE endpoints are password or token protected to prevent malicious actors from accessing user data and code.
@@
The AI4EOSC Platform allows executing federated learning servers [64], allowing AI/ML models to be trained collaboratively across multiple institutions without sharing private data. The platform currently supports the two most widely used and robust frameworks: Flower [10] and NVFLARE [76]. Moreover, we have extended Flower to support token-based client-server authentication [78], with secret management directly implemented via the dashboard. We have also implemented novel aggregation methods, such as FedAvgOpt [79], and incorporated server-side metric differential privacy [77] with fixed clipping, in order to prevent attackers from retrieving training data at inference time from the aggregated model and to defend against client inference attacks.
In order to participate in the federated training, federated clients can be deployed within the AI4EOSC Platform or connect from outside, thus easily accommodating the individual privacy needs of the different data owners. This flexible and distributed approach for training models allows leveraging a wide range of resources in the Cloud continuum that could not support a full ML training cycle by themselves due to their low capacity (e.g., edge devices).
3.1.5 Continuous Integration and Delivery of models
A key component of the AI4EOSC platform, that underpins the MLOps model management is the CI/CD system comprising both platform-enforced and user defined pipelines executed via Jenkins [46]. This system ensures that any AI module integrated in the Marketplace satisfies the same consistent requirements: e.g. by ensuring that published models include the needed metadata, implementing data and model drift detection, or triggering updates on existing services based on a given module.
@@ 3.1.5 Continuous Integration and Delivery of models
    Metadata validation: We ensure that the user-defined metadata (4.1) is compliant with the metadata schema that the platform relies on. This is a critical step, taking into account that the PAPI, Dashboard, and the interoperability with other metadata schemes require that accurate and consistent metadata is in place.
@@
    Docker images generation and publication: The platform relies on Docker containers to ensure that the modules are self-contained and executable in a wide variety of systems. Therefore we trigger the build of new images and their publication in the platform registry [^3] https://registry.cloud.ai4eosc.eu/ and Docker Hub [^4] https://hub.docker.com/u/ai4oshub.
@@ 3.1.6 MLOps components
Users deploying AI modules (3.1.3) or Development Environments (3.1.4) can log their training metrics into a dedicated MLflow [93] tracking server instance we have deployed, featuring a custom Authentication GUI to be able to integrate it with our authentication system. MLflow credentials, saved in the platform secrets manager and will then be automatically injected in the WMS deployments (3.1.1) created from the Dashboard (3.1.2).
@@
The AI4EOSC consortium has developed a drift monitoring system, DriftWatch [^5] https://drift-watch.cloud.ai4eosc.eu/. This component allows that whenever AI modules (3.1.3) are created, users can choose to fit a detector on labeled clean and faulty data. Then, at inference time, when the models are deployed in production, the detector can be used to collect and send metrics to the DriftWatch server. From an intuitive UI, users are able to log in and visualize whether drift has occurred, eventually inspecting the data that caused the drift.
In order to detect if a drift is happening or not, a detector is needed. Driftwatch is agnostic to the underlying drift detection framework chosen by the developer such as Alibi-detect [89], Evidently [35], etc. However, AI4EOSC has also developed Frouros [20], a Python library for drift detection in machine learning systems that supports a large combination of classical and more recent algorithms for both concept and data drift detection. Frouros has been integrated within the platform, in order to provide drift detection within AI pipelines easily [86].
3.1.7 Provenance
Our system is designed to keep track of all the steps involved in building of a module: from data provenance, to model training, hardware characteristics, etc. Once a module has committed changes to the Git repository, the CI/CD pipeline triggers the fetching of the provenance metadata, gathering information from multiple sources: metadata (4.1) and training resources (3.1.1) from the Platform API (3.1.2), experiment tracking metrics from MLflow (3.1.6) and build information from the CI/CD pipeline itself (3.1.5).
These non-semantic and non-structured metadata fragments are stored in JSON format in a PostgreSQL database [87] managed by an API built with Spring [12]. Then, using the RDF Mapping Language (RML) via CARML [17], we convert the plain JSON metadata into an RDF provenance graph (W3C PROV [90]), optionally using domain specific ontologies. Having the provenance data into RDF allows for an easier conversion to external formats like FAIR4ML [19] or any additional mapping. Moreover, this approach allows us to extend the provenance information with virtually any information that the platform or the user is able to provide to the system.
Figure 6: Example of a provenance graph built for a demo application.
To present the metadata to the users, we provide visual tools integrated in the dashboard (cf. Figure 6), as well as integration with a chatbot assistant (Section 3.3), allowing to perform semantic queries against the graph using natural language.
@@ 3.2 AI/ML as a Service system
The AI as a Service (AIaaS), shown in Figure 7, allows user to deploy AI modules (3.1.3) in a serverless inference platform, based on OSCAR [71, 40].
Figure 7: AI4EOSC AI/ML as a Service architecture: container view (C4 notation).
The OSCAR cluster allows to autoscale resources on demand, automatically deploying multiple instances of the model to run predictions as needed. The platform allows both synchronous and asynchronous calls to the endpoints via file uploads to a MinIO object store [60].
The AI4EOSC Dashboard (3.1.2) allows users to automatically create and manage those serverless endpoints for any AI module in the catalog. Thanks to the DEEPaaS API [37] that interfaces the module, users are able to send inference requests to any AI module in a standardized way.
@@
Pipelines can currently be created with two different visual composition tools: Node-Red [65] whose instances are managed via the OSCAR cluster, and Elyra [29], integrated with EGI Jupyter Notebooks. Users can choose between these two tools to compose their pipelines based on their working environment needs and prior experience.
@@ 3.3 Large Language Models Services system
This system, described in Figure 8, is designed to provide a federated inference system focused on the needs of Large Language Models (LLMs). The system leverages vLLM [53] and LiteLLM [55] in order to provide an easy way to include additional LLM providers, that can be geographically distributed. Moreover, it is integrated with the platform secret management system, in order to be able to manage API keys for the users. This system can be directly used via its OpenAI compatible API endpoints, in order to be integrated with any user tool, or via a conversational engine. Moreover, this system is key to provide AI4EOSC platform users with an specific support agent, able to support users with direct answers from the documentation knowledge base.
Figure 8: AI4EOSC Large Language Model Services architecture: container view (C4 notation).
@@ 3.4 Platform orchestration system
This system, depicted in Figure 9 is in charge of assisting resource providers and platform operators by automatically provisioning the resources needed to deploy the rest of the platform services. In the AI4EOSC platform, we use the INDIGO-PaaS Orchestration system [26, 80] to handle the deployment of platform components across the different IaaS platforms made available by the federated Cloud providers.
Figure 9: AI4EOSC platform orchestration architecture: container view (C4 notation).
@@
    The Federation-registry [81], which is the configuration management database for dynamic cloud federation. Built using FastAPI [73] and neo4j [62], the Federation-registry is able to efficiently manage the federation-related information and SLA agreements.
@@
    The INDIGO-PaaS Orchestrator Dashboard, which provides a user-friendly interface for managing and monitoring the TOSCA-based deployments, built on Flask [68] and MySQL [61].
@@
4 Other platform design considerations
4.1 AI4EOSC metadata
In the current data-driven ecosystem, rich metadata is of paramount importance in order to comprehensively describe a given asset in an homogeneous way. Well defined metadata allows also to work towards interoperability, both at the platform and at the AI/ML model level. To ensure that all the AI modules developed in the platform provide consistent metadata, the AI4EOSC Platform enforces that each AI module (3.1.3) or tool (3.1.4) that is registered in the AI4EOSC catalog must include a valid metadata file following a defined schema implemented via JSON Schema [47].

### 05 May 2026 v3

@@ 3 The AI4EOSC platform architecture
In the following section, we describe the architecture of AI4EOSC using a simplified view of the C4 Model [11] to offer a clear and structured visualization of its components and interactions. This simplified representation aims to enhance the article readability. For a detailed and comprehensive view of the AI4EOSC architecture, please refer to the complete C4 implementation available at [27].
Our architecture follows a modular design, organized into four distinct systems: AI4EOSC Development Platform, AI4EOSC AI/ML as a Service, AI4EOSC LLM services, and AI4EOSC Platform Orchestration. These components work interconnected in order to be able to deliver an specific functionality, but being integrated with the rest of the systems. This modularity is illustrated in Figure 1. The legend for all the C4 diagrams is shown in Figure 2.
Figure 1: Simplified view of the AI4EOSC architecture system landscape view (C4 notation).
@@ 3.1 AI4EOSC Development Platform system
Figure 3: Simplified AI4EOSC Development platform architecture: container view (C4 notation).
@@ 3.1.1 Workload Management System
The Workload Management System (WMS) is a federation of service providers that is used to manage the platform resources and execute the platform jobs. Given the federated nature of the platform, we have chosen a set of components that allow us to federate geographically distributed resources easily in the form of a service mesh. Therefore, in order to address the challenges and complexity of cross-datacenter resource federation, the AI4EOSC architecture leverages a decentralized WMS based on Hashicorp Nomad and Consul. Unlike centralized orchestrators, like Kubernetes, this combination allows for a lightweight scheduling across geographically distributed providers with minimal administrative overhead. Moreover, both the inter-cluster orchestration and intra-cluster scheduling is managed from the same control plane, without the need of additional layers or components. The current federation includes resources from over 5 different clouds located in 5 different countries: IFCA-CSIC (Spain), IISAS (Slovakia), INCD (Portugal), Tubitak (Turkey) and PNSC (Poland). On top of Nomad node failure monitoring, the AI4EOSC platform runs periodic quality checks (including configuration and performance) to remove low performing nodes from the federation.
The AI4EOSC WMS handles two primary types of executions: user-oriented workloads and system-level management tasks. User workloads refer to both AI modules (3.1.3) and other specialized tools (3.1.4) that the platform offers to enrich the user experince. Both of these can be executed in two distinct modes. On the one hand, in standard mode resources are persistent, providing researchers with continuous access via Interactive Development Environments (IDEs) or APIs. On the other hand, the batch mode utilizes a transient allocation strategy, where resources are provisioned strictly for the duration of a training process to maximize global cluster efficiency and resource sharing.
Besides, in order to enhance user experience and provide structural functionality to the platform, AI4EOSC augments Nomad jobs with transparent sidecar tasks. These components manage additional function that is not directly exposed the users, such as automated connectivity to external storage systems, perform dataset synchronization from remote repositories at runtime, and trigger user notifications. Moreover, system-level jobs including Traefik reverse proxies, the Platform API, and the administrative Dashboard (3.1.2) are also part of these workloads to maintain platform stability and service accessibility without direct user intervention.
@@ 3.1.2 Platform API and dashboard
Built on top of the API, the AI4EOSC Dashboard is the main user entry-point to the AI4EOSC Platform. Both PAPI and the dashboard are built with extensibility in mind, in order to be able to easily integrate new functionality without interfering with the current one.
@@
The Dashboard (Figure 4) allows users to easily explore and view the associated metadata for the different catalogs available in the system: AI modules (3.1.3), tools (3.1.4), LLM services (3.3) and other external AI catalogs (4.2.2). It also provides the ability to launch development environments or codespaces in the platform, batch executions, easy management of existing deployments, platform-wide secret management (e.g., API keys, storage integration) and resource consumption reporting, among other user-facing functionalities.
@@ 3.1.3 AI Modules
The AI modules are the main platform asset that users interact with, as they encapsulate the AI/ML models that the platform is able to manage. They are both produced and consumed by the users, although the AI4EOSC development team also maintains reference implementations of relevant AI models. They are created and stored as git repositories (currently hosted on GitHub, but any other software forge can be used) allowing users to easily manage and upload new updates to the modules, as well as making it possible to enforce quality assurance mechanisms and MLOps practices (3.1.5) from the platform point of view. To support the development of new AI modules that comply with the AI4EOSC platform and best practices (e.g., metadata standards (4.1.1)), a set of pre-defined software templates is provided to users.
@@ 3.1.4 Tools
For technical reasons, some of these tools do not make sense to be deployed per-user and therefore are instead offered as platform-wide services, like the experiment tracking service or the drift monitoring (3.1.5); however the integration mechanism remains the same.
@@
The AI4EOSC Platform allows users to deploy a ready-to-use development environment, referred to as Codespaces, to create their new AI modules (3.1.3). The environment can be used out-of-the-box with multiple versions of the main Deep Learning frameworks such as TensorFlow [2], PyTorch [44]. Additional frameworks are also supported by starting from provided plain Ubuntu or NVIDIA CUDA flavors, or by installing a custom software stack.
When configuring the deployment, users can choose their preferred IDE for code development. The platform currently supports both Visual Studio Code (VSCode) bundled with many useful extensions, and installed using Code Server, and JupyterLab. All IDE endpoints are password or token protected to prevent malicious actors from accessing user data and code.
@@
The AI4EOSC Platform allows executing federated learning servers [40], allowing AI/ML models to be trained collaboratively across multiple institutions without sharing private data. The platform currently supports the two most widely used and robust frameworks: Flower [9] and NVFLARE [47]. Both FL servers can be launched interactively from the dashboard, and FL clients can be deployed within the AI4EOSC Platform or connect from outside, thus easily accommodating the individual privacy needs of the different data owners. This flexible and distributed approach allows leveraging a wide range of resources in the Cloud continuum that could not support a full ML training cycle by themselves due to their low capacity (e.g., edge devices).
Beyond the standard framework capabilities, AI4EOSC has extended the Flower framework with a set of platform-level enhancements. First, we have implemented token-based client-server authentication [50], managed directly from the dashboard via the platform secret management system. This allows operators to create and revoke client tokens at any time, ensuring that only authorized clients can participate in the training. Second, the server incorporates a mechanism to monitor the divergence between the model weights of different clients, allowing the detection of clients that deviate significantly from the rest and may be introducing poisoning behavior, with the option to revoke their access token immediately. Third, server-side differential privacy can be applied during aggregation using fixed-clipping Gaussian mechanisms, and the platform further incorporates the notion of server-side metric differential privacy [48], which scales the noise multiplier dynamically based on a distance metric computed from local model parameters, protecting against client inference attacks. Finally, AI4EOSC introduces FedAvgOpt [51], a novel aggregation method that optimizes the aggregated weights to be as close as possible to the individual client weights using the Nelder-Mead method, improving convergence without compromising privacy. The supported aggregation strategies (i.e., FedAvg, FedAvgM, FedMedian, FedProx, FedOpt, FedAdam, FedYogi, and FedAvgOpt) are selectable directly from the dashboard.
The security and privacy implications of this learning approach, including its role in defending against poisoning and inference attacks, are further discussed in Section 4.4.3.
3.1.5 MLOps practices
Continuous Integration and Delivery of models
A key component of the AI4EOSC platform, that underpins the MLOps model management is the CI/CD system comprising both platform-enforced and user defined pipelines executed via Jenkins. This system ensures that any AI module integrated in the Marketplace satisfies the same consistent requirements: e.g. by ensuring that published models include the needed metadata, or triggering updates on existing services based on a given module.
@@ 3.1.5 MLOps practices
    Metadata validation: We ensure that the user-defined metadata (4.1.1) is compliant with the metadata schema that the platform relies on. This is a critical step, taking into account that the PAPI, Dashboard, and the interoperability with other metadata schemes require that accurate and consistent metadata is in place.
@@
    Docker images generation and publication: The platform relies on Docker containers to ensure that the modules are self-contained and executable in a wide variety of systems. Therefore we trigger the build of new images and their publication in the platform Harbor container registry [^2] https://registry.cloud.ai4eosc.eu/ and Docker Hub [^3] https://hub.docker.com/u/ai4oshub. To avoid unnecessary builds, the rebuild step is only triggered when non-metadata changes are detected and can be disabled in specific git branches.
@@
Users deploying AI modules (3.1.3) or Development Environments (3.1.4) can log their training metrics into a dedicated MLflow [60] tracking server instance we have deployed, featuring a custom Authentication GUI to be able to integrate it with our authentication system. MLflow credentials are saved in the platform secrets manager and automatically injected in the WMS deployments (3.1.1) created from the Dashboard (3.1.2).
@@
AI4EOSC has developed a drift monitoring system, DriftWatch [^4] https://drift-watch.cloud.ai4eosc.eu/. This component allows to monitor AI modules (3.1.3) in production, warning the model developer when performance starts to degrade and a model retraining is needed. In order to detect if a drift is happening or not, a detector is needed. Driftwatch is agnostic to the underlying drift detection framework chosen by the developer such as Alibi-detect [57], Evidently, etc. However, AI4EOSC has developed Frouros [17], a Python library for drift detection in machine learning systems that supports a large combination of classical and more recent algorithms for both concept and data drift detection. When an AI module is developed, users can define reference clean/faulty data and train a Frouros detector on them (users handling image data would need to first train an autoencoder to reduce the dimensionality of their data). Then, at inference time, when the models are deployed in the production AIaaS (3.2), the inference data is passed trough the Frouros detector and the resulting p-values for detection are sent to the DriftWatch server [18]. From an intuitive user interface, users are able to log in and visualize whether drift has occurred, eventually inspecting the data that caused the drift.
@@ 3.2 AI/ML as a Service system
The AIaaS, shown in Figure 6, allows user to deploy AI modules (3.1.3) in a serverless inference platform, based on OSCAR [45].
Figure 6: Simplified AI4EOSC AI/ML as a Service architecture: container view (C4 notation).
The OSCAR cluster allows to autoscale resources on demand, automatically deploying multiple instances of the model to run predictions as needed. The platform allows both synchronous and asynchronous calls to the endpoints via file uploads to a MinIO object store. The AI4EOSC Dashboard (3.1.2) allows users to automatically create and manage those serverless endpoints for any AI module in the catalog. Thanks to the DEEPaaS API [28] that interfaces the module, users are able to send inference requests to any AI module in a standardized way.
@@
Pipelines can currently be created with two different visual composition tools: Node-Red, whose instances are managed via the OSCAR cluster, and Elyra, integrated with Jupyter-based environments like the AI4EOSC Development Environment (3.1.4) or the EGI Notebooks. Users can choose between these two tools to compose their pipelines based on their working environment needs and prior experience.
@@ 3.3 Large Language Models Services system
This system, described in Figure 7, is designed to provide a federated inference system focused on the needs of Large Language Models (LLMs). The system leverages vLLM [37] and LiteLLM in order to provide an easy way to include additional LLM providers, that can be geographically distributed. To achieve this, each resource provider deploys their own LLMs with vLLM and aggregates them with a local LiteLLM instance. Then, in turn, the platform’s LiteLLM instance aggregates the different local LiteLLMs, effectively serving all provider LLMs under a single endpoint.
Moreover, the AI4EOSC Dashboard allows users to easily manage API keys to access the service. This system can be directly used via its OpenAI compatible API endpoints, in order to be integrated with any user tool, and via the platform’s conversational engine based on OpenWebUI. Moreover, this system is key to provide AI4EOSC platform users with an specific support agent, able to support users with direct answers from the documentation knowledge base.
Figure 7: Simplified AI4EOSC Large Language Model Services architecture: container view (C4 notation).
@@ 3.4 Platform orchestration system
This system, depicted in Figure 8, is in charge of assisting resource providers and platform operators by automatically provisioning the resources needed to deploy the rest of the platform services. In the AI4EOSC platform, we use the INDIGO-PaaS Orchestration system [23, 52] to handle the deployment of platform components across the different IaaS platforms made available by the federated Cloud providers.
Figure 8: Simplified AI4EOSC platform orchestration architecture: container view (C4 notation).
@@
    The Federation-registry [53], which is the configuration management database for dynamic cloud federation. Built using FastAPI and neo4j, the Federation-registry is able to efficiently manage the federation-related information and SLA agreements.
@@
    The INDIGO-PaaS Orchestrator Dashboard, which provides a user-friendly interface for managing and monitoring the TOSCA-based deployments, built on Flask and MySQL.
@@
4 Platform design considerations
4.1 Platform FAIR design
As mentioned when presenting the motivation, the FAIR principles are central to the design of the AI4EOSC platform. Ensuring that AI modules and their associated assets are transparently discoverable, well-described, and reusable by the broader scientific community is a core requirement for a research-oriented platform. AI4EOSC addresses this through standardized metadata schemas, comprehensive end-to-end provenance tracking, and a commitment to openness. This approach ensures that every asset produced by the platform can be understood, reproduced, and built upon by others.
4.1.1 AI4EOSC metadata
In the current data-driven ecosystem, rich metadata is of paramount importance in order to comprehensively describe a given asset in an homogeneous way. Well defined metadata allows also to work towards interoperability, both at the platform and at the AI/ML model level. To ensure that all the AI modules developed in the platform provide consistent metadata, the AI4EOSC Platform enforces that each AI module (3.1.3) or tool (3.1.4) that is registered in the AI4EOSC catalog must include a valid metadata file following a defined schema implemented via JSON Schema.